# CGCOD: Class-Guided Camouflaged Object Detection

### v1

@@ Abstract
Camouflaged Object Detection (COD) is designed to identify objects that blend seamlessly with their surroundings. Due to the complexity of camouflaged objects (such as shape, color, and texture), their semantic cues are often blurred or completely lost, posing a significant challenge for COD. Existing COD methods often rely on visual features, which are not stable enough in changeable camouflage environments. This instability leads to false positives and false negatives, resulting in incomplete or inaccurate segmentation results. In this paper, to solve this problem, we propose a new task, Class-Guided Camouflaged Object Detection (CGCOD), which extends the traditional COD task by introducing object class knowledge, significantly improving the robustness and segmentation accuracy of the model in complex environments. Toward this end, we construct a dataset, CamoClass, containing the camouflaged objects in the real scenes and their corresponding class annotation. Based on this, we propose a multi-stage framework CGNet which consists of a plug-and-play class prompt generator and a class-guided detector. Under the guidance of textual information, CGNet enables efficient segmentation. It is worth emphasizing that for the first time, we extend the object class annotations on existing COD benchmark datasets, and introduce a flexible framework to improve the performance of the existing COD model under text guidance.
@@ 1 Introduction
These challenges arise from relying solely on purely visual paradigm, which fail to capture the full semantic characteristics of camouflaged objects. This limitation becomes particularly pronounced in scenes where objects closely resemble their backgrounds or are heavily occluded. Camouflaged objects exhibit strong visual similarity to their background (e.g., shape, color, scale), yet maintain significant semantic differences However, existing benchmark datasets primarily focus on visual information and lack stable, explicit features to effectively emphasize these semantic differences. As a result, purely visual models often suffer from high rates of false positives (e.g., misclassifying background regions as objects) and false negatives (e.g., failing to detect actual objects), leading to a substantial decline in detection accuracy. Moreover, various factors such as environmental lighting, interference, and occlusion, can profoundly affect the visual appearance of the object, making it challenging for models to consistently extract distinctive object features. For instance, partial occlusion distorts the original structure and contour of the object, reducing the model’s ability to recover complete semantic information from limited visual cues.
@@
In recent years, integrating visual and textual features has proven effective for enhancing image content understanding, thereby improving performance in various multi-modal tasks Tan and Bansal (2019); Lu et al. (2019). Targeted textual prompts provide explicit semantic cues that help models focus on object-related features and reduce background interference Radford et al. (2021). However, existing COD benchmark datasets provide only visual information, while current visual-textual COD methods Zhang et al. (2024) depend on coarse-grained text generated by large language models. Therefore, it is of great significance for the COD community to build a dataset that integrates both visual and textual information and develop a new COD framework.
@@
Motivated by the above discussion, we introduce a new task, Class-Guided Camouflaged Object Detection (CGCOD). This task accurately focus on relevant regions that likely contain camouflaged objects by leveraging the specific class text, thereby improving model’s detection accuracy and robustness. To support our research, we construct a new dataset, CamoClass. It integrates multiple COD datasets, including CAMO Le et al. (2019), COD10K Fan et al. (2022), CHAMELEON Przemysław et al. (2018), and NC4K Lv et al. (2021). We annotate all images with their corresponding class labels. To further explore the role of class guidance in COD, we propose a new method CGNet, which consist a plug-and-play class prompt generator (CPG) and a novel high-performance baseline class-guided detector (CGD). The CGD is designed to seamlessly integrate with CPG. In particular, we introduce a Class Semantic Guidance (CSG) consisting of two multi-head self-attention (MHSA Vaswani (2017)) modules as a bridge between CPG and CGD. The CSG aims to fuse information from multiple sources and reduce domain differences between them. Fig. 1 shows the ability of CGNet for the CGCOD task. In the CPG, we introduce a cross-modal multi-head attention module to enhance class semantics by facilitating interactions between visual and textual features. Additionally, our multi-level visual collaboration module (MVCM) ensures comprehensive alignment and integration of these features at various depths, enabling cooperative extraction of object semantics. Together, these modules generate class prompt features that guide the COD detector for precise segmentation. In the CGD, we design a semantic consistency module (SCM) to reduce false negatives and positives by leveraging the class prompt features from the CPG. CGNet further explores multi-modal feature correlations, suppresses background noise, and delivers accurate predictions.
@@
  • We propose a new task, CGCOD, which is the first time to fully explore the performance of COD model under class guidance and provide a new insight for traditional COD approaches.
  • We construct a novel dataset CamoClass by integrating existing COD datasets and annotate the class description of the camouflaged object.
  • We design CGNet, a new method featuring a plug-and-play class prompt generator that enhances existing COD models, and a high-performance baseline class-guided detector optimized for the CPG.
  • Comprehensive experiments demonstrate the superior performance of CGNet, which not only outperforms existing state-of-the-art methods, but also incorporates the CPG plug-and-play module to enhance the performance of other detectors.
@@ 2 Related Works
Referring Image Segmentation (RIS). RIS is a fundamental computer vision task aiming to segment object regions in an image based on natural language descriptions. Early RIS methods are often built upon CNN-LSTM architectures Hu et al. (2016). They encode image and text information separately, and leverage simple attention mechanisms. This limits cross-modal interaction, making it challenging for models to selectively focus on object-relevant features, resulting in coarse segmentation. For example, Wang et al. (2022) introduces a vision-language decoder that maps fine-grained text semantics to pixel-level representations for improved modality alignment. Tang et al. (2023) employs a mask classification framework with token-based querying to enhance object-level understanding and cross-modal reasoning. Yue et al. (2024) uses a word-guided dual-branch structure to dynamically align visual features with linguistic cues, boosting RIS performance. Although our method also integrates textual and visual modalities, our objective differs substantially from traditional RIS. RIS tasks use natural language to segment clearly described objects, whereas our task combines object class knowledge to detect camoflaged objects blending into the background, making it more challenging. This fundamental difference in goals and application distinguishes our work from RIS research.
@@
Camouflaged Object Detection (COD). In nature, animals use camouflage as a defense mechanism to evade predators Copeland and Trivedi (1997). The high complexity of camouflaged strategies makes COD a consistently challenging task. Early COD methods relied on handcrafted features like texture Galun et al. (2003), gradient Zhang et al. (2016), and color Pulla Rao et al. (2020) to distinguish foreground from background but suffered from limited feature extraction capabilities, especially in complex scenes. In contrast, deep learning-based approaches, such as CNNs and Transformers, automatically learn camouflaged object features, offering significant improvements. For example, Fan et al. (2020a) uses densely connected features and receptive fields to mimic a predator’s search and recognition process. Pang et al. (2022) integrates multi-scale features and hierarchical units to emulate human visual patterns when observing blurry images, capturing hybrid-scale semantics. Sun et al. (2022) leverages edge cues to guide the model in perceiving object structures. Song et al. (2023) introduces a non-local token enhancement mechanism and a feature compression decoder to enhance feature interactions. He et al. (2023) employs a learnable wavelet transform to decompose features into different frequency domains, effectively distinguishing foreground from background. With the emergence of large models, Yu et al. (2024) utilizes SAM’s zero-shot capability by combining depth and RGB features for precise segmentation. Additionally, Zhang et al. (2024); Ruan et al. (2024) use environment descriptions generated by large language models to investigate how background camouflage attributes affect object detection. These approaches demonstrate deep learning’s flexibility and innovation in the COD task, offering diverse solutions.
Despite many advances, existing COD methods rely too heavily on visual features, which fail to adequately address the complex characteristics of camouflaged objects. This limitation may not only lead to poor model robustness, but also cause misrecognition problems.
In contrast, our work introduces class-level textual information into the COD domain, along with a new benchmark dataset.
By explicitly incorporating class semantics, our model can more accurately focus on the object region in the image and effectively reduce the interference of the complex environment on object perception.
@@ 3 Proposed CamoClass Dataset
In order to advance the study of visual-textual approaches in COD, we propose a new dataset, CamoClass. The focus of this research is to improve object detection in camouflage scenes by integrating the category knowledge of objects. This method extends the traditional idea of COD and provides a new benchmark. It should be emphasized that our innovation lies not only in the class annotation itself, but also in how to use class annotation to provide COD with new perspective and idea in visual-textual interaction without changing the existing model training paradigm.
Annotation Strategy. CamoClass contains 8,663 samples from 128 class, of which 44 are multi-class samples. Although the number of images of each class is uneven (as shown in Fig. 2), reflecting the varying frequencies of different camouflaged objects in the real world, this also poses challenges for the model’s generalization and robustness. Each sample consists of an image ($\mathbf{I}$), ground truth ($\mathbf{GT}$), and edge information ($\mathbf{E}$). In the process of labeling, we adopt strict class definition and standardization strategy to reduce ambiguity and ensure clarity and consistency of class definition, and unify the treatment of fine-grained class such as “small fish” and “black cat”. Grouping them into broader class (such as “fish” and “cat”) simplifies the class space and makes it easier to identify object. For images containing multiple-class objects, we provide an independent segmentation mask for each class. For example, an image containing both a "dog" and a "cat" will have separate masks for each class, ensuring accurate and high-quality annotation.
Flexible Application. CamoClass is highly flexible and extensible. Researchers can continue to use the training and test partitioning of the original COD dataset to maintain comparability with previous studies and ensure that the introduction of new information does not affect existing experimental paradigms. In addition, we integrate these sub-datasets to form a large COD dataset with diverse classes, and further divide the test set into seen classes $C_{seen}$ (present in the training set) and unseen classes $C_{unseen}$ (not present in the training set). The integrated dataset provides sufficient training data for the class-guided COD task, supporting the joint semantic modeling of text and image, and the location of camouflaged object in complex scenes. It is expected that this large-scale dataset will become a cornerstone of community research, driving the development of the COD field in open-vocabulary Pang et al. (2025), referring Ji et al. (2024), and zero-shot learning Ren et al. (2023).
@@ 4 Proposed Method
When students are in class, they don’t just rely on memorization, but externalize information by recording key concepts or points of knowledge. This focus on clearly targeted content helps them to grasp the learning points in a targeted way. Inspired by this principle, we introduce the CGNet to alleviate the limitations faced by the purely visual COD model. Fig. 3 provides an overview of our network architecture. In particular, the input of the CGNet is composed of tow parts, which include image and class-level textual information of the camouflaged object, termed as $I_{r}$ and $I_{t}$. They also comes from the dataset CamoClass, which are fed into the class prompt generator (CPG) and class-guided detector (CGD) to generate class-prompt feature $G_{c}$ and the segmentation result $P_{1}$.
@@ 4.1 Class Prompt Generator (CPG)
Overview. The CPG is designed to integrate both visual and textual information to generate prompt features for object classes. By leveraging textual information, it enhances the semantic understanding of the object’s context in the image, which helps mitigate the impact of visually similar but irrelevant objects. This mechanism significantly reduces both false positives and false negatives, thus improving the precision of the COD task. Specifically, in this phase, we first employ a frozen CLIP textual encoder to generate textual embeddings $F_{t}$ from the class-level textual information. Simultaneously, we extract multi-level visual features $\{F_{i}\}_{i=1}^{3}$ from 8, 16, and 24 layers of the frozen CLIP visual encoder. These features are then fused via a FPN module Wang et al. (2022) and decode through a Transformer block to generate preliminary multi-modal features $F_{m}$. To strengthen the alignment between text and pixels and emphasize class semantics of object, we introduce a cross-modal multi-head attention (CMA). The computation process follows the basic attention mechanism, projecting features from both modalities into different embedding spaces, and calculating the correlation between text and image features across multiple subspaces. This highlights the class semantic knowledge of object in the image, ultimately generating the feature $F_{c}$. Furthermore, considering the effectiveness of multi-level feature interactions in image understanding, we propose a Multi-level Visual Collaboration module (MVCM) to enrich visual features at different levels and generate the feature $F_{v}$. Fig. 4 visualizes the stages within the CPG, demonstrating progressive improvements in object localization and validating the effectiveness of our approach.
@@
Multi-level Visual Collaboration Module (MVCM). Existing research has shown that the collaboration between low and high level structural information is essential for fine-grained segmentation, reflecting principles akin to human visual processing. In order to take advantage of the coarse segmentation multi-level features of the CLIP visual encoder to generate more fine-grained visual features, we propose the MVCM to enhance the multi-layer features $F_{c}$, $F_{2}$ and $F_{3}$. Specially, this module consists of two stages: the alignment stage and the enhancement stage. In the alignment stage, we employ two Multi-Head Self-Attention (MHSA) modules to align $F_{c}$ with $F_{2}$ and $F_{c}$ with $F_{3}$, followed by a summation operation to fuse them, generating intermediate features $F_{\text{n}}$. These features are then further enhanced.
$F_{\text{n}}=\text{MHSA}(F_{c},F_{2})\oplus\text{MHSA}(F_{c},F_{3})$ (1)
@@
In the enhancement stage, we apply a residual-based progressive refinement strategy on $F_{\text{n}}$. This strategy progressively strengthens the feature, ultimately producing the fused feature representation $F_{v}$. This design aims to maximize the retention of fine details while extracting rich semantic information, thereby improving the model’s ability to detect camouflaged objects. which are formulated as:
$F_{n}^{1}=\text{Conv}(\text{Conv}(F_{n})\oplus F_{n})$ (2)
$F_{n}^{2}=\text{Conv}(\text{Conv}(F_{n}\otimes F_{n}^{1})\oplus F_{n}\otimes F_{n}^{1})$ (3)
$F_{n}^{3}=\text{Conv}(\text{Conv}(F_{n}\otimes F_{n}^{2})\oplus F_{n}\otimes F_{n}^{2})$ (4)
$F_{v}=\text{Conv}(\text{Concat}(F_{n}^{1},F_{n}^{2},F_{n}^{3})).$ (5)
Here, Conv() signifies a $3\times 3$ convolution.

### v2

@@ Abstract
Camouflaged Object Detection (COD) aims to identify objects that blend seamlessly into their surroundings. The inherent visual complexity of camouflaged objects, including their low contrast with the background, diverse textures, and subtle appearance variations, often obscures semantic cues, making accurate segmentation highly challenging. Existing methods primarily rely on visual features, which are insufficient to handle the variability and intricacy of camouflaged objects, leading to unstable object perception and ambiguous segmentation results. To tackle these limitations, we introduce a novel task, class-guided camouflaged object detection (CGCOD), which extends traditional COD task by incorporating object-specific class knowledge to enhance detection robustness and accuracy. To facilitate this task, we present a new dataset, CamoClass, comprising real-world camouflaged objects with class annotations. Furthermore, we propose a multi-stage framework, CGNet, which incorporates a plug-and-play class prompt generator and a simple yet effective class-guided detector. This establishes a new paradigm for COD, bridging the gap between contextual understanding and class-guided detection. Extensive experimental results demonstrate the effectiveness of our flexible framework in improving the performance of proposed and existing detectors by leveraging class-level textual information. Code: https://github.com/bbdjj/CGCOD.
@@ 1 Introduction
These challenges arise from relying solely on purely visual paradigm, which fail to capture the full semantic characteristics of camouflaged objects. This limitation becomes particularly pronounced in scenes where objects closely resemble their backgrounds or are heavily occluded. Camouflaged objects exhibit strong visual similarity to their background (e.g., shape, color, scale), yet maintain significant semantic differences. However, existing benchmark datasets primarily focus on visual information and lack stable, explicit information to effectively emphasize these semantic differences. As a result, purely visual models often suffer from high rates of false positives (e.g., misclassifying background regions as objects) and false negatives (e.g., failing to detect actual objects), leading to a substantial decline in detection accuracy. Moreover, various factors such as environmental lighting, interference, and occlusion, can profoundly affect the visual appearance of the object, making it challenging for models to consistently extract distinctive object features. For instance, partial occlusion distorts the original structure and contour of the object, reducing the model’s ability to recover complete semantic information from limited visual cues.
@@
In recent years, integrating visual and textual features has proven effective for enhancing image content understanding, thereby improving performance in various multi-modal tasks Tan and Bansal (2019). Targeted textual prompts provide explicit semantic cues that help models focus on object-related features and reduce background interference Radford et al. (2021). However, existing COD benchmark datasets provide only visual information, while current visual-textual COD methods Zhang et al. (2025) depend on coarse-grained text generated by large language models. Therefore, it is of great significance for the COD community to build a dataset that integrates both visual and textual information and develop a new COD framework.
@@
Motivated by the above discussion, we introduce a new task, Class-Guided Camouflaged Object Detection (CGCOD). This task accurately focus on relevant regions that likely contain camouflaged objects by leveraging the specific class text, thereby improving model’s detection accuracy and robustness. To support our research, we construct a new dataset, CamoClass. It integrates multiple COD benchmark datasets, including CAMO Le et al. (2019), COD10K Fan et al. (2022), CHAMELEON Przemysław et al. (2018), and NC4K Lv et al. (2021). We annotate all images with their corresponding class information. To further explore the role of class guidance in COD, we propose a new method, CGNet, which consist a plug-and-play class prompt generator (CPG) and a novel high-performance baseline class-guided detector (CGD). The CGD is designed to seamlessly integrate with CPG. And we introduce a class semantic guidance as a bridge between CPG and CGD to fuse information from multiple sources and reduce domain differences between them. Fig. 1 shows the ability of CGNet for the CGCOD task.
In the CPG, we introduce a cross-modal multi-head attention module to enhance class semantics by facilitating interactions between visual and textual features. Additionally, our multi-level visual collaboration module (MVCM) ensures comprehensive alignment and integration of these features at various levels, enabling cooperative extraction of object semantics. Together, these modules generate class prompt feature $G_{c}$ that guide the COD detector for precise segmentation. In the CGD, we design a semantic consistency module (SCM) to reduce false negatives and positives by leveraging the class prompt features from the CPG. CGNet further explores multi-modal feature correlations, suppresses background noise, and achieves accurate predictions.
@@
  • We propose a new task, CGCOD, which is the first time to fully explore the performance of COD methods under the class guidance and provide a new insight for the traditional COD field.
  • We construct a novel dataset CamoClass by integrating existing COD datasets and annotate the class label of the camouflaged object.
  • We design CGNet, a new method featuring a plug-and-play class prompt generator (CPG) that enhances existing COD detectors, and a high-performance baseline class-guided detector optimized for the CPG.
  • Extensive experiments demonstrate the superior performance of our CGNetover existing state-of-the-art methods. Further, the CPG’s contribution to existing detectors’ performance demonstrates its strong pluggability.
@@ 2 Related Works
Referring Image Segmentation (RIS). RIS is a fundamental computer vision task aiming to segment object regions in an image based on natural language descriptions. Early RIS methods are often built upon CNN-LSTM architectures Hu et al. (2016). They encode image and text information separately, and leverage simple attention mechanisms. This limits cross-modal interaction, making it challenging for models to selectively focus on object-relevant features, resulting in coarse segmentation. For example, Wang et al. (2022) introduces a vision-language decoder that maps fine-grained text semantics to pixel-level representations for improved modality alignment. Tang et al. (2023) employs a mask classification framework with token-based querying to enhance object-level understanding and cross-modal reasoning. Yue et al. (2024) uses a word-guided dual-branch structure to dynamically align visual features with linguistic cues, boosting RIS performance. Although our method also integrates textual and visual modalities, our objective differs substantially from traditional RIS. RIS tasks use natural language to segment clearly described objects, whereas our task combines object class knowledge to detect camoflaged objects blending into the background, making it more challenging.
@@
Camouflaged Object Detection (COD). In nature, animals use camouflage as a defense mechanism to evade predators Copeland and Trivedi (1997). The high complexity of camouflaged strategies makes COD a consistently challenging task. Early COD methods relied on handcrafted features like texture Galun et al. (2003), gradient Zhang et al. (2016), and color Pulla Rao et al. (2020) to distinguish foreground from background but suffered from limited feature extraction capabilities, especially in complex scenes. In contrast, deep learning-based approaches, such as CNNs and Transformers, automatically learn camouflaged object features, offering significant improvements. For example, Fan et al. (2020a) uses densely connected features and receptive fields to mimic a predator’s search and recognition process. Pang et al. (2022) integrates multi-scale features and hierarchical units to emulate human visual patterns when observing blurry images, capturing hybrid-scale semantics. Sun et al. (2022) leverages edge cues to guide the model in perceiving object structures. Song et al. (2023) introduces a non-local token enhancement mechanism and a feature compression decoder to enhance feature interactions. He et al. (2023) employs a learnable wavelet transform to decompose features into different frequency domains, effectively distinguishing foreground from background. With the emergence of large models, Yu et al. (2024) utilizes SAM’s zero-shot capability by combining depth and RGB features for precise segmentation. Additionally, Zhang et al. (2025); Ruan et al. (2024) use environment descriptions generated by large language models to investigate how background camouflage attributes affect object detection. These approaches demonstrate deep learning’s flexibility and innovation in the COD task, offering diverse solutions. Despite many advances, existing COD methods rely too heavily on visual features, which fail to adequately address the complex characteristics of camouflaged objects. This limitation may not only lead to poor model robustness, but also cause misrecognition problems. In contrast, our work introduces class-level textual information into the COD domain, along with a new benchmark dataset. By explicitly incorporating class semantics, our model can more accurately focus on the object region in the image and effectively reduce the interference of the complex environment on object perception.
@@ 3 Proposed CamoClass Dataset
In order to advance the study of visual-textual approaches in COD, we propose a new dataset, CamoClass. The focus of this research is to improve object detection in camouflaged scenes by integrating the class-level textual information of objects. This method extends the traditional paradigm of COD and provides a new benchmark. It should be emphasized that our innovation lies not only in the class annotation itself, but also in how to use class annotation to provide COD with new perspective and idea in visual-textual interaction without changing the existing COD model training paradigm.
Annotation Strategy. CamoClass contains 10,523 samples from 128 class, of which 44 are multi-class samples. Each sample consists of an image (${I}$), ground truth (${G}$), and edge information (${E}$). The number of images in each class is uneven (as illustrated in Fig. 2), which reflects the real-world frequency of different camouflaged objects, but this imbalance also poses challenges for the model’s generalization and robustness. In the process of labeling, we adopt strict class definition and standardization strategy to reduce ambiguity and ensure clarity and consistency of class definition, and unify the treatment of fine-grained class such as “small fish” and “black cat”. Grouping them into broader class (such as “fish” and “cat”) simplifies the class space and makes it easier to identify object. For images containing multiple-class objects, we provide an independent segmentation mask for each class. For example, an image containing both a "dog" and a "cat" will have separate masks for each class, ensuring accurate and high-quality annotation.
Flexible Application. CamoClass is highly flexible and extensible. Researchers can continue to use the training and test partitioning of the original COD dataset to maintain comparability with previous studies and ensure that the introduction of new information does not affect existing experimental paradigms. In addition, we integrate these sub-datasets to form a large camouflaged object dataset with class information, and further divide the test set into seen classes $\mathcal{C}_{seen}$ (present in the training set) and $\mathcal{C}_{unseen}$ (not present in the training set) to test the generalization of COD model on $\mathcal{C}_{seen}$ classes. The detailed criteria for this division can be found in the supplementary material. CamoClass dataset provides sufficient training data for the CGCOD task, supporting the joint semantic modeling of text and image, and the location of camouflaged object in complex scenes. It is expected that this large-scale dataset will become a cornerstone of community research, driving the development of the COD field in open-vocabulary Pang et al. (2025), referring Ji et al. (2024), and zero-shot learning Ren et al. (2023).
@@ 4 Proposed Method
When students learn in a classroom, they do not merely rely on rote memorization but actively externalize information by recording key concepts. This focus on clearly targeted content helps them to grasp the learning points in a targeted way. Inspired by this, we introduce the CGNet to alleviate the limitations faced by the purely visual COD model. As illustrated in Fig. 3, the input of the CGNet is composed of tow parts, which include image and class-level textual information of the camouflaged object, termed as $I_{r}$ and $I_{t}$. They come from the dataset CamoClass, which are fed into the class prompt generator, class-semantic guidance and class-guided detector to generate class-semantic prompt feature $G_{c}$ and the segmentation result $P_{1}$. The class-semantic guidance consists of two MHSA modules to reduce domain difference among feathers from different sources.
@@ 4.1 Class Prompt Generator (CPG)
Overview. The CPG is designed to integrate both visual and textual information to generate prompt feature for COD detector. Integrating textual information allow the model to capture deeper contextual semantics of objects and effectively reduce interference from visually similar yet semantically irrelevant features. This mechanism significantly suppresses both false positives and false negatives, thereby improving the precision of the COD task. Specifically, in this phase, we first employ a frozen CLIP textual encoder to generate textual embeddings $F_{t}$ from the class-level textual information. Simultaneously, we extract multi-level visual features $\{F_{i}\}_{i=1}^{3}$ from 8, 16, and 24 layers of the frozen CLIP visual encoder. These features are then fused via a FPN module Wang et al. (2022) and decode through a Transformer block to generate preliminary multi-modal features $F_{m}$. To further strengthen the alignment between text and pixels and emphasize class semantic features of object, we introduce a cross-modal multi-head attention (CMA). The computation process follows the basic attention mechanism, projecting features from both modalities into different embedding spaces, and calculating the correlation between text and image features across multiple subspaces. This highlights the class semantic knowledge of object in the image, ultimately generating the feature $F_{c}$. Furthermore, considering the effectiveness of multi-level feature collaboration in image understanding, we propose a multi-level visual collaboration module (MVCM) to enrich visual features at different levels and generate the deep multimodal feature $F_{v}$. Fig. 4 visualizes the stages within the CPG, demonstrating progressive improvements in object localization and validating its effectiveness.
@@
Multi-level Visual Collaboration Module (MVCM). Existing research has shown that the collaboration between low and high level structural information is essential for fine-grained segmentation, reflecting principles akin to human visual processing. In order to take advantage of the coarse segmentation multi-level features of the CLIP visual encoder to generate more fine-grained visual features, we propose the MVCM to enhance the multi-layer features $F_{c}$, $F_{2}$ and $F_{3}$, where $F_{c}$ represents low-level multimodal features, and $F_{2}$ and $F_{3}$ correspond to high-level visual features.
To better understand and integrate information from different sources, this module consists of two stages: the alignment stage and the enhancement stage. In the alignment stage, we employ two multi-Head self-attention (MHSA) modules to align $F_{c}$ with $F_{2}$ and $F_{c}$ with $F_{3}$, followed by a summation operation to fuse them, generating intermediate features $F_{n}$.
$F_{{n}}={MHSA}(F_{c},F_{2})+{MHSA}(F_{c},F_{3})$ (1)
@@
In the enhancement stage, we apply a residual-based progressive refinement strategy on $F_{{n}}$. This strategy progressively strengthens the feature, ultimately producing the deep multimodal fused feature representation $F_{v}$. This design aims to maximize the retention of fine details while extracting rich semantic information, thereby improving the model’s ability to detect camouflaged object. which are formulated as:
$F_{n}^{1}={Conv_{3}}({Conv_{3}}(F_{n})+F_{n})$ (2)
$F_{n}^{2}={Conv_{3}}({Conv_{3}}(F_{n}\otimes F_{n}^{1})+F_{n}\otimes F_{n}^{1})$ (3)
$F_{n}^{3}={Conv_{3}}({Conv_{3}}(F_{n}\otimes F_{n}^{2})+F_{n}\otimes F_{n}^{2})$ (4)
$F_{v}={Conv_{3}}({Concat}(F_{n}^{1},F_{n}^{2},F_{n}^{3})).$ (5)
where $\otimes$, $Conv_{3}$, and $Concat$ denote Hadamard multiplication, a 3 × 3 convolutional layer, and the concatenation operation, respectively.